Completely and thoroughly explain how a Proposed Income Statement and a Proposed Return on Investment (ROI) could be used to make a decision on whether or not to add a new product to the existing product line of a company that makes frozen food products.
Answer #1

Let us consider the following.

There is company called Org that has a new proposal to produce product A.

Here are the results from product A

Existing A Total
Sales $       80,000 $ 45,000 $1,25,000
COGS $       56,000 $ 35,000 $    91,000
GP / Contribution $       24,000 $ 10,000 $    34,000
Fixed Costs $       10,000 $ 11,000 $    21,000
Net Income $       14,000 $ -1,000 $    13,000

If these were the actual results, Product A would have been scrapped since the total company income has reduced.

This is how a Proposed Income Statement can help in making decision regarding addition of a new product.

Now suppose the total assets (investment) was $ 50,000. Another $ 15,000 were required for the new product.

Existing ROI = 14,000 / 50,000 = 28%
Proposed ROI = 13,000 / 65,000 = 20%

Using a Proposed Income Statement helps us to know the impact of the new product on the total income of the company. If the total income after considering the new product is less than the original income, then the new product must not be started. However, if the situation is other way around, that is, the total income is more with the addition of the new product, the product must be started.

Same is the case with ROI. ROI give us the rate at which income was generated. It tells us what portion of the investment was earned during a period. If the proposed ROI is higher with the new product, it must be started.